OpenStudy (anonymous):

Please help with geometry! In a given circle, the radius is 48 cm. find the measure of the circles diameter.

OpenStudy (anonymous):

d = r/2

OpenStudy (anonymous):

r*2=D

OpenStudy (anonymous):

Would that be 24 cm?

OpenStudy (anonymous):

d= 96cm

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

r=48 r*2=D So 96

OpenStudy (anonymous):

yay I'm correct!

OpenStudy (anonymous):

Your supposed to explain

OpenStudy (anonymous):

I'm dumb

OpenStudy (anonymous):

r= d/2 I meant, derp.

OpenStudy (anonymous):

lol